OpenStudy (anonymous):

Evaluate : tgx= 3/4 ; sin^3 2x- cos^3 2x / 2+4sinx

OpenStudy (dmb1215):

@Mertsj we need some help

OpenStudy (mertsj):

what does it mean tgx=3/4

OpenStudy (anonymous):

tangent X= 3/4

OpenStudy (mertsj):

The abbreviation for tangent is tan

OpenStudy (anonymous):

sorry(

OpenStudy (anonymous):

any idea how to solve it?

OpenStudy (mertsj):

\[\frac{\sin ^3(2x)-\cos ^3(2x)}{2+4\sin x}=\frac{(\sin (2x)-\cos (2x))(\sin ^2(2x)+\sin (2x)\cos (2x)+\cos ^2(2x)}{2+4\sin x}\] \[\frac{(\sin (2x)-\cos (2x))(1+\sin (2x)\cos (2x))}{2+4\sin x}=\]

OpenStudy (mertsj):

\[\frac{(2\sin xcos x-(1-2\sin ^2x))(1+(2\sin x \cos x)(1-2\sin ^2x))}{2+4\sin x}=\]

OpenStudy (mertsj):

\[\frac{(2\sin x \cos-1+2 \sin ^2x)(1+2\sin x \cos x-4\sin ^3x \cos x)}{2+4\sin x}\]

OpenStudy (mertsj):

If tan x = 3/4 then sin x = 3/5 and cos x = 4/5 so you can substitute and finish it.

OpenStudy (anonymous):

okay!) Mertsj, thank you a lot!!!) that's gonna be on my exam tmrw)

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

You might be able to make it a bit easier by saying: sin(2x)=2sinxcosx=2(3/5)(4/5)=24/25 And cos(2x)=1-2sin^2x=1-2(3/5)^2=1-2(9/25)=1-18/25=7/25

OpenStudy (mertsj):

Then you could substitute into this step: \[\frac{(\sin (2x)-\cos (2x))(1+\sin (2x)\cos (2x))}{2+4\sin x}\]